OpenStudy (molly):

Explain how to solve the equation 3t + 11 = -10 show that the solution works.

OpenStudy (aroub):

3t+11=-10 Add -11 to both sides: 3t=-21 divide 3 to both sides: t=-7 check: susbtitute t by -7 in the original equation: 3(-7)+11=-10 -21+11=-10 -21=-21

OpenStudy (anonymous):

You beat me to it . Haha .

OpenStudy (molly):

lol thankss but im putting another question up so check it outt